# Anytime-valid t-tests and confidence sequences for Gaussian means with unknown variance

### v4

@@ 1 Introduction
Sequential extensions of the t-test were studied by Rushton [1950], Hoel [1954], Ghosh [1960], Sacks [1965], who were interested in calculating the approximate thresholds to control the type 1 error at a particular predefined stopping time. Inspired by the sequential Z-test work of Robbins, Lai [1976] constructed a sequential t-test that controls a different type of error that will be of particular interest to us: nonasymptotic, conservative (i.e. non-approximate) time-uniform control of the type 1 error at all arbitrary stopping times, possibly not specified in advance. We will define this formally later, but we mention at the outset that the filtration will matter very much: some methods will have guarantees at any stopping times in the original data filtration, while others only have guarantees in a reduced filtration, and studying this phenomenon in the context of t-tests will be an important part of this paper.
Lai and more recent authors [Grünwald et al., 2023, Pérez-Ortiz et al., 2022] situate the Gaussian t-test problem into the broader framework of group-invariant tests; indeed, the t-test null is invariant under rescaling of observations by a constant. The tests are mostly constructed using variants of the sequential likelihood ratio test (LRT) by Wald [1945], utilizing the fact that the likelihood ratios are nonnegative martingales (see Section 2.5). Notably, Pérez-Ortiz et al. [2022] shows the group-invariant extension of the LRT to be optimal (in the “e-power” sense to be defined in Section 2.3) in the orbit-versus-orbit ($\mu/\sigma=\theta_{0}$ versus $\mu/\sigma=\theta_{1}$) setting. However, their results fail to generalize to the composite alternative $\mu/\sigma\neq\theta_{0}$. Building upon these group invariance tools, Lindon et al. [2022] construct a sequential F-test for the regression set-up, which coincides with one of the t-tests that we shall present when no covariate exists. A slightly different approach is taken by Jourdan et al. [2023] who study the same problem in the bandit context, sequentially bounding the empirical mean and variance separately using existing Z-test-type bounds and obtaining via a union bound a sequence of confidence “boxes” for mean and variance simultaneously. We shall return to and compare our bounds with some of these results later.
We shall focus in this paper primarily on the construction of e-processes, which are sequences of test statistics whose realized values reflect the evidence against the null in favor of the alternative. These e-processes can be thresholded and/or inverted to yield nonasymptotic, conservative, “anytime-valid” hypothesis tests and confidence intervals (called “confidence sequences”), but we emphasize that they are worthy of being used in their own right as well. Indeed, these e-processes directly measure evidence against the null: values below one are expected under the null, and large(r) values indicate (more) evidence against the null.
@@
    We apply the method of “Universal Inference”, a variant of the LRT by Wasserman et al. [2020], to construct a t-test e-process for the point null $\mu=\mu_{0}$, as well as for the one-sided composite null $\mu\leqslant\mu_{0}$, implying respectively a two-sided and a one-sided confidence sequence for the mean $\mu$, both of which are new.
@@
    We fill in the missing details of Lai’s [1976] sequential t-test, looking into the curious question of why his approach yields a confidence sequence without an e-process, using the theory of non-$L^{1}$ nonnegative supermartingales [Wang and Ramdas, 2023a].
@@
    We derive the information-theoretic lower bound on the width of confidence intervals, and the upper bound on the growth of e-processes for the t-test, both of which are attained, or almost attained, by our results (Contributions 1 and 3 in this list). We also show that other conservative sequential t-tests available in the literature (Lai’s original method, the plug-in method by Jourdan et al. [2023], the median method by Howard and Ramdas [2022]), and very surprisingly, even the fixed-time classical t-test, fail to be optimal in these regards.
@@
    We conduct a detailed comparison among tests we derive and in the literature in terms of (a) which filtration they are safe for optional stopping in (the larger the better); (b) what their e-powers are against a fixed alternative (the larger the better); and (c) for their derived confidence sequences, the width or expected width, especially as a function of sample size $n$ and level $\alpha$ (smaller is better).
@@ 2.1 Notations
A universal probability space $(\Omega,\mathcal{A},\mathbb{P})$ is used when we consider the randomness of data. We use the indexed letters $X_{1},X_{2},\dots$ to denote an infinite stream of observations, taking values in $\mathbb{R}$. We denote by $\{\mathcal{F}_{n}\}_{n\geqslant 0}$ the canonical filtration generated by the data, i.e. $\mathcal{F}_{n}=\sigma(X_{1},\dots,X_{n})$. Our shorthand notations on the sample include the partial sum $S_{n}=\sum_{i=1}^{n}X_{i}$, the sample average $\widehat{\mu}_{n}=S_{n}/n$, the partial sum of squares $V_{n}=\sum_{i=1}^{n}X_{i}^{2}$, the average square $\overline{X^{2}_{n}}=V_{n}/n$, the sample variance without Bessel’s correction $v_{n}=V_{n}/n-S_{n}^{2}/n^{2}$, and the t-statistic $T_{n-1}=\sqrt{n-1}\frac{S_{n}-n\mu_{0}}{\sqrt{nV_{n}-S_{n}^{2}}}$ where $\mu_{0}$ is the actual mean of $X_{1},X_{2},\dots$.
Usually, we consider i.i.d. observations $X_{1},X_{2},\dots$; and we use italic $P$ etc. to denote distributions over $\mathbb{R}$, while calligraphic $\mathcal{P}$ etc. to denote classes of distributions. For sample-dependent events and random variables, we carefully adopt the notations $P(\,\cdot\,)$ as
$\mathbb{P}[\,\cdot\,]\quad\text{when }X_{1},X_{2},\dots\overset{\mathrm{iid}}{\sim}P,$ (1)
and $\mathrm{E}_{P}(\,\cdot\,)$ as
$\mathbb{E}[\,\cdot\,]\quad\text{when }X_{1},X_{2},\dots\overset{\mathrm{iid}}{\sim}P.$ (2)
The conditional expectation $\mathrm{E}_{P}(\,\cdot\,|\mathcal{F})$ given $\mathcal{F}\subseteq\mathcal{A}$ is defined similarly. Univariate normal distribution with mean $\mu$ and variance $\sigma^{2}$ will be referred to as $N_{\mu,\sigma^{2}}$ – so “$N_{0,1}(X_{1}>0)=0.5$” means “$\mathbb{P}[X_{1}>0]=0.5$ when $X_{1},X_{2},\dots\overset{\mathrm{iid}}{\sim}$ standard normal”; its density is denoted by $p_{\mu,\sigma^{2}}$. The class of all normal distributions on $\mathbb{R}$ (excluding the degenerate $\sigma=0$ case) is denoted by $\mathcal{N}$. Classes of normal distributions with specified mean (or range of means) but any positive variance are denoted by, e.g., $\mathcal{N}_{\mu=\mu_{0}}$, $\mathcal{N}_{\mu\leqslant\mu_{0}}$; those with specified standardized mean by, e.g., $\mathcal{N}_{\mu/\sigma=\theta_{0}}$.
@@ 2.2 Sequential Statistics
The classical sequential analysis traces back to Wald [1947], and is formulated as testing the null $H_{0}:P\in\mathcal{P}_{0}$ against the alternative $H_{1}:P\in\mathcal{P}_{1}$, making a decision (to reject or not) each time a new data point $X_{n}\sim P$ is seen until a rejection is made. The type 1 error $P(H_{0}\text{ is ever rejected})$ for $P\in\mathcal{P}_{0}$ is to be controlled under a prescribed $\alpha$. This leads to the definition of any-time valid p-value [Johari et al., 2022], as a stochastic process $\{p_{n}\}_{n\geqslant 1}$ such that $P(p_{\tau}\leqslant p)\leqslant p$ for any stopping time $\tau$, any $p\in(0,1)$, and any $P\in\mathcal{P}_{0}$. Reject at the hitting time $\inf\{n:p_{n}\leqslant\alpha\}$ clearly controls the type I error.
@@
The classical duality between tests and interval estimates also manifests in the sequential setting. Let $\mathcal{P}$ be a family of distributions and let $\theta:\mathcal{P}\to\mathbb{R}$ be the parameter of interest. A $(1-\alpha)$-confidence sequence (CS) [Darling and Robbins, 1967] over $\mathcal{P}$ for $\theta$ is a sequence of confidence intervals $\{\operatorname{CI}_{n}\}_{n\geqslant 1}$ such that, for any $P\in\mathcal{P}$, $P(\forall n\geqslant 1,\;\theta(P)\in\operatorname{CI}_{n})\geqslant 1-\alpha$; or equivalently, $P(\theta(P)\in\operatorname{CI}_{\tau})\geqslant 1-\alpha$ for any stopping time $\tau$. If the null $\mathcal{P}_{0}\subseteq\mathcal{P}$ is defined as a preimage $\theta^{-1}(\theta_{0})$, a sequential test can be constructed via rejection $H_{0}$ whenever $\theta_{0}\notin\operatorname{CI}_{n}$, and vice versa.
In this paper, the asymptotic tightness of a confidence sequence is measured in terms of (1) its growth rate, the rate at which its width increases when $n$ and the sample are fixed while $\alpha\to 0$ (or while $\alpha\downarrow\alpha_{\star}(n)$ if $\alpha$ is not allowed to be below $\alpha_{\star}(n)$); and (2) its shrinkage rate, the rate at which its width decreases when $\alpha$ is fixed while $n\to\infty$). For example, it is known that CSs for the mean over all 1-subGaussian random variables, and over all random variables with variance $\leqslant 1$ can both attain the minimax rates of growth $\sqrt{\log(1/\alpha)}$ and shrinkage $\sqrt{\log\log(n)/n}$ [Howard et al., 2021, Waudby-Smith and Ramdas, 2023, Wang and Ramdas, 2023b].
2.3 Test processes and their e-power
@@
Sequential tests and confidence sequences are usually constructed via nonnegative supermartingales (NSMs) and more generally, e-processes [Howard et al., 2021, Ramdas et al., 2022]. Let us, following Ruf et al. [2023], define them and state their key properties in composite terms.
@@
Recalling that $\mathcal{P}$ denotes a class of distributions, if, for all $P\in\mathcal{P}$, the process $\{M_{n}\}$ is NSM (or NM, ENSM) for $P$, we say it is an NSM (or NM, ENSM) for $\mathcal{P}$.
Further, we say $\{M_{n}\}$ is
  • an e-process for $P$ on $\{\mathcal{G}_{n}\}$, if $M_{n}$ is $\mathcal{G}_{n}$-measurable and, for each $P\in\mathcal{P}$, there is an NSM $\{N_{n}^{P}\}$ with $N_{0}^{P}=1$ on $\{\mathcal{G}_{n}\}$ for $P$ that upper bounds $\{M_{n}\}$: $P(M_{n}\leqslant N_{n}^{P})=1$.
In particular, when only $n=1$ observation is considered, we call a nonnegative random variable $M=M(X_{1})$ with $\sup_{P\in\mathcal{P}}\mathrm{E}_{P}(M)\leqslant 1$ an e-value for $\mathcal{P}$.
Note that the filtration $\{\mathcal{G}_{n}\}$ is usually the canonical one $\{\mathcal{F}_{n}\}$, but as we shall see later, other non-trivial choices will lead to interesting results.
@@
Lemmas 2.1 and 2.2 suggest that, if $\mathcal{P}_{0}$ is a set of null distributions, NSMs, e-processes, and ENSMs for $\mathcal{P}_{0}$ can be seen as measurements of evidence against the null $H_{0}:P\in\mathcal{P}_{0}$. To see that, if we have at hand an NSM or e-process $\{M_{n}\}$ for $\mathcal{P}_{0}$ such that $M_{0}=1$, we may reject the null $\mathcal{P}_{0}$ whenever the process $M_{n}$ exceeds $1/\alpha$. This, due to Lemma 2.1, controls the sequential type 1 error $P(H_{0}\text{ is ever rejected})$ within $\alpha$ for any $P\in\mathcal{P}_{0}$. Similarly, if we have an ENSM $\{M_{n}\}$ for $\mathcal{P}_{0}$, we may set the right hand side of (5) to $\alpha$ and solve the corresponding $\varepsilon$, rejecting $H_{0}:P\in\mathcal{P}_{0}$ whenver $M_{n}$ exceeds $\varepsilon$. We refer to NSMs, e-processes, and ENSMs for $\mathcal{P}_{0}$ as test processes for the null $\mathcal{P}_{0}$. Desirably, a powerful test process shrinks under the null $\mathcal{P}_{0}$, and grows under the alternative $\mathcal{P}_{1}$ so rejection can happen. To measure the power of a test process under a fixed alternative distribution $Q\in\mathcal{P}_{1}$, we define the e-power of a test process $\{M_{n}\}$ as the $Q$-almost sure limit $\liminf_{n\to\infty}\frac{\log M_{n}}{n}$; for a single e-value $M=M(X_{1})$, its e-power is defined as the expected value $\mathrm{E}_{Q}(\log M)$ [Vovk and Wang, 2024].
A confidence sequence for a parameter $\theta$ follows in the same way if we have an NSM or an ENSM $\{M^{\theta_{0}}_{n}\}$ for every $\mathcal{P}_{\theta_{0}}=\{P\in\mathcal{P}:\theta(P)=\theta_{0}\}$: by solving the inequality $M_{n}^{\theta_{0}}\leqslant 1/\alpha$ or $M_{n}^{\theta_{0}}\leqslant\varepsilon$ in $\theta_{0}$, that is. An illustrative comparison between NSMs and ENSMs in the Gaussian case with known variance for both constructing tests and confidence sequences is provided recently by Wang and Ramdas [2023a, Section 5].

### v5

@@ 1 Introduction
Sequential extensions of the t-test were studied by Rushton [1950], Hoel [1954], Ghosh [1960], Sacks [1965], who were interested in calculating the approximate thresholds to control the type 1 error at a particular predefined stopping time. Inspired by the sequential Z-test work of Robbins, Lai [1976] constructed a sequential t-test that controls a different type of error that will be of particular interest to us: nonasymptotic, conservative (i.e. non-approximate) time-uniform control of the type 1 error at all arbitrary stopping times, possibly not specified in advance. We will define this formally later, but we mention at the outset that the filtration will become an interesting matter as different sequential methods are defined on different filtrations. Studying this phenomenon in the context of t-tests will be an important part of this paper.
Lai and more recent authors [Grünwald et al., 2024, Pérez-Ortiz et al., 2024] situate the Gaussian t-test problem into the broader framework of group-invariant tests; indeed, the t-test null is invariant under rescaling of observations by a constant.
The tests are mostly constructed using variants of the sequential likelihood ratio statistic by Wald [1945], utilizing the fact that the likelihood ratios are nonnegative martingales (see Section 2.6). Notably, Pérez-Ortiz et al. [2024] shows the group-invariant extension of the likelihood ratio statistic to be optimal (in the “e-power” sense to be defined in Section 2.3) primarily in the point-versus-point ($\mu/\sigma=\theta_{0}$ versus $\mu/\sigma=\theta_{1}$) setting. We shall closely discuss their contributions later. Building upon these group invariance tools, Lindon et al. [2022] construct a sequential F-test for the regression set-up, which coincides with one of the t-tests that we shall present when no covariate exists. A slightly different approach is taken by Jourdan et al. [2023] who study the same problem in the bandit context, sequentially bounding the empirical mean and variance separately using existing Z-test-type bounds and obtaining via a union bound a sequence of confidence “boxes” for mean and variance simultaneously. We shall return to and compare our bounds with some of these results later.
We shall focus in this paper primarily on the construction of e-processes, which are sequences of test statistics whose realized values reflect the evidence against the null in favor of the alternative. These e-processes can be thresholded and/or inverted to yield nonasymptotic, conservative, “anytime-valid” hypothesis tests and confidence intervals (called “confidence sequences”), but we emphasize that they have clear interpretations without thresholding them to yield tests. Indeed, these e-processes directly measure evidence against the null: values below one are expected under the null, and large(r) values indicate (more) evidence against the null.
@@
    We apply the method of “Universal Inference”, a variant of the likelihood ratio test by Wasserman et al. [2020], to construct a t-test e-process for the point null [^2] In this paper, a “point null” refers to a null hypothesis that a parameter of interest takes a specific value. Since it possibly contains multiple distributions, this is different from a “simple null”. $\mu=\mu_{0}$, as well as for the one-sided composite null $\mu\leqslant\mu_{0}$, implying respectively a two-sided and a one-sided confidence sequence for the mean $\mu$, both of which are new.
@@
    We fill in the missing details of Lai’s [1976] sequential t-test, looking into the curious question of why his approach yields a confidence sequence without an e-process, using the theory of non-integrable nonnegative supermartingales [Wang and Ramdas, 2023a].
@@
    We derive the information-theoretic lower bound on the width of confidence intervals, and the upper bound on the growth of e-processes for the t-test, both of which are attained, or almost attained, by our results (Contributions 1 and 3 in this list). We also show that other conservative sequential t-tests available in the literature (Lai’s original method, the plug-in method by Jourdan et al. [2023], the median method by Howard and Ramdas [2022]), and very surprisingly, even the fixed-time classical t-test, fail to be optimal in the rate with respect to the sample size $n$ and/or the type 1 error level $\alpha$.
@@
    We conduct a detailed comparison among tests we derive and in the literature in terms of (a) which filtration they are safe for optional stopping in (the larger the better); (b) what their e-powers are against a fixed alternative (the larger the better); and (c) for their derived confidence sequences, the width or expected width, especially as a function of sample size $n$ and level $\alpha$.
@@ 2.1 Notations
A universal probability space $(\Omega,\mathcal{A},\mathbb{P})$ is used when we consider the randomness of data, where the expectation is denoted by $\mathbb{E}$. We use the indexed letters $X_{1},X_{2},\dots$ to denote an infinite stream of observations, taking values in $\mathbb{R}$. We denote by $\{\mathcal{F}_{n}\}_{n\geqslant 0}$ the canonical filtration generated by the data, i.e. $\mathcal{F}_{n}=\sigma(X_{1},\dots,X_{n})$. Our shorthand notations on the sample include the partial sum $S_{n}=\sum_{i=1}^{n}X_{i}$, the sample average $\widehat{\mu}_{n}=S_{n}/n$, the partial sum of squares $V_{n}=\sum_{i=1}^{n}X_{i}^{2}$, the average square $\overline{X^{2}_{n}}=V_{n}/n$, the sample variance without Bessel’s correction $s_{n}^{2}=V_{n}/n-S_{n}^{2}/n^{2}$, and the t-statistic $T_{n-1}=\sqrt{n-1}\frac{S_{n}-n\mu_{0}}{\sqrt{nV_{n}-S_{n}^{2}}}$ where $\mu_{0}$ is the actual mean of $X_{1},X_{2},\dots$.
Usually, we consider i.i.d. observations $X_{1},X_{2},\dots$; and we use italic $P$ etc. to denote distributions over $\mathbb{R}$, while calligraphic $\mathcal{P}$ etc. to denote classes of distributions. For a sample-dependent event $A$, we use the notation $P(A)$ for the probability $\mathbb{P}[A]$ when $X_{1},X_{2},\dots\overset{\mathrm{iid}}{\sim}P$; and for a random variable realized as a function of the sample $Z=Z(X_{1},\dots,X_{n})$, we write $\mathrm{E}_{P}(Z)$ for the expectation $\mathbb{E}[Z]$ when $X_{1},X_{2},\dots\overset{\mathrm{iid}}{\sim}P$. The conditional expectation $\mathrm{E}_{P}(\,\cdot\,|\mathcal{F})$ given $\mathcal{F}\subseteq\mathcal{A}$ is defined similarly. Univariate normal distribution with mean $\mu$ and variance $\sigma^{2}$ will be referred to as $N_{\mu,\sigma^{2}}$ – so “$N_{0,1}(X_{1}>0)=0.5$” means “$\mathbb{P}[X_{1}>0]=0.5$ when $X_{1},X_{2},\dots\overset{\mathrm{iid}}{\sim}$ standard normal”; its density is denoted by $p_{\mu,\sigma^{2}}$. The class of all normal distributions on $\mathbb{R}$ (excluding the degenerate $\sigma=0$ case) is denoted by $\mathcal{N}$. Classes of normal distributions with specified mean (or range of means) but any positive variance are denoted by, e.g., $\mathcal{N}_{\mu=\mu_{0}}$, $\mathcal{N}_{\mu\leqslant\mu_{0}}$; those with specified standardized mean by, e.g., $\mathcal{N}_{\mu/\sigma=\theta_{0}}$.
@@ 2.2 Sequential Statistics
The theory of sequential hypothesis testing is formulated as testing the null $H_{0}:P\in\mathcal{P}_{0}$ against the alternative $H_{1}:P\in\mathcal{P}_{1}$, where the statistician makes a decision (to reject, accept, or sample further) each time a new data point $X_{n}\sim P$ is seen until rejection or acceptance. In the formulation of Wald [1945, 1947], the decision scheme is required to control the type 1 error rate $P(H_{0}\text{ is ever rejected})$ for $P\in\mathcal{P}_{0}$ within a prescribed level $\alpha$, and the type 2 error rate $P(H_{0}\text{ is ever accepted})$ for $P\in\mathcal{P}_{1}$ within $\beta$. A slightly different formulation, proposed by Darling and Robbins [1968], only requires the type 1 error rate to be controlled within $\alpha$, but the decision schemes under this constraint are often proved to have “power one”, i.e., $P(H_{0}\text{ is ever rejected})=1$ for $P\in\mathcal{P}_{1}$. We shall follow the latter formulation. We derive sequential tests of type 1 error rate strictly bounded by $\alpha$ (“level-$\alpha$ sequential tests”) and prove their power guarantees. These power guarantees are stated in terms of “e-power”, to be formally introduced in Section 2.4, which is a finer-grained quantification of power than the rejection rate $P(H_{0}\text{ is ever rejected})$ under $P\in\mathcal{P}_{1}$.
@@
The classical duality between tests and interval estimates also manifests in the sequential setting. Let $\mathcal{P}$ be a family of distributions and let $\theta:\mathcal{P}\to\mathbb{R}$ be the parameter of interest. A $(1-\alpha)$-confidence sequence (CS) [Darling and Robbins, 1967] over $\mathcal{P}$ for $\theta$ is a sequence of confidence intervals $\{\operatorname{CI}_{n}\}_{n\geqslant 1}$ such that, for any $P\in\mathcal{P}$, $P(\forall n\geqslant 1,\;\theta(P)\in\operatorname{CI}_{n})\geqslant 1-\alpha$; or equivalently, $P(\theta(P)\in\operatorname{CI}_{\tau})\geqslant 1-\alpha$ for any stopping time $\tau$ (see Lemma 3 of Howard et al. [2021]). If the null $\mathcal{P}_{0}\subseteq\mathcal{P}$ is defined as a preimage $\theta^{-1}(\theta_{0})$, a sequential test can be constructed via rejection $H_{0}$ whenever $\theta_{0}\notin\operatorname{CI}_{n}$, and vice versa. In this paper, the asymptotic tightness of a confidence sequence is measured in terms of (1) its growth rate, the rate at which its width increases when $n$ and the sample are fixed while $\alpha\to 0$ (or while $\alpha\downarrow\alpha_{\star}(n)$ if $\alpha$ is not allowed to be below $\alpha_{\star}(n)$); and (2) its shrinkage rate, the rate at which its width decreases when $\alpha$ is fixed while $n\to\infty$). For example, it is known that the minimax rates of growth $\sqrt{\log(1/\alpha)}$ and shrinkage $\sqrt{\log\log(n)/n}$ can be obtained in CSs for the mean over all 1-subGaussian random variables [Howard et al., 2021, Waudby-Smith and Ramdas, 2023], and also all random variables with variance upper bounded by some constant $\sigma^{2}$ [Wang and Ramdas, 2023b]. As we shall see soon, the optimal rates for the t-test are larger.
2.3 Test Processes and Their Maximal Inequalities
@@
There has been a recent revival of the idea, originally due to Robbins [1970], of constructing sequential tests and confidence sequences via nonnegative supermartingales (NSMs) and more generally, e-processes [Howard et al., 2021, Ramdas et al., 2022]. Let us define them below.
@@
We say the process $\{M_{n}\}$ is NSM (or NM, ENSM) for $\mathcal{P}$ if it is an NSM (or NM, ENSM) for all $P\in\mathcal{P}$. When an NSM (or NM) for $\mathcal{P}$ has an initial expected value $\mathrm{E}_{P}(M_{0})=1$ for all $P\in\mathcal{P}$, we call it a test supermartingale (or test martingale) for $\mathcal{P}$. We call a nonnegative random variable $M=M(X_{1},\dots,X_{n})$ with $\sup_{P\in\mathcal{P}}\mathrm{E}_{P}(M)\leqslant 1$ an e-value for $\mathcal{P}$. Further, we say the process $\{M_{n}\}$ is
  • an e-process for $\mathcal{P}$ on $\{\mathcal{G}_{n}\}$ if for any $\{\mathcal{G}_{n}\}$-adapted stopping time $\tau$, $M_{\tau}$ is an e-value for $\mathcal{P}$; i.e., $\sup_{P\in\mathcal{P}}\mathrm{E}_{P}(M_{\tau})\leqslant 1$.
Note that the filtration $\{\mathcal{G}_{n}\}$ is usually the canonical one $\{\mathcal{F}_{n}\}$, but as we shall see later, other non-trivial choices will lead to interesting results. NMs and NSMs for $\mathcal{P}$ are clearly e-processes for $\mathcal{P}$, on the same filtration. More generally, the following equivalent characterization of e-processes is due to Ramdas et al. [2020, Lemma 6]:
  • A nonnegative process $\{M_{n}\}$ adapted to $\{\mathcal{G}_{n}\}$ is an e-process for $\mathcal{P}$ on $\{\mathcal{G}_{n}\}$ if and only if for each $P\in\mathcal{P}$, there is a test supermartingale $\{N_{n}^{P}\}$ for $P$ on $\{\mathcal{G}_{n}\}$ that upper bounds $\{M_{n}\}$: $P(M_{n}\leqslant N_{n}^{P})=1$.
@@
Lemmas 2.1 and 2.2 suggest that, if $\mathcal{P}_{0}$ is a set of null distributions, NSMs, e-processes, and ENSMs for $\mathcal{P}_{0}$ can be seen as measurements of evidence against the null $H_{0}:P\in\mathcal{P}_{0}$. To see that, if we have at hand an NSM or e-process $\{M_{n}\}$ for $\mathcal{P}_{0}$ such that $M_{0}=1$, we may reject the null $\mathcal{P}_{0}$ whenever the process $M_{n}$ exceeds $1/\alpha$. This, due to Lemma 2.1, controls the sequential type 1 error $P(H_{0}\text{ is ever rejected})$ within $\alpha$ for any $P\in\mathcal{P}_{0}$. Similarly, if we have an ENSM $\{M_{n}\}$ for $\mathcal{P}_{0}$, we may set the right hand side of (3) to $\alpha$ and solve the corresponding $\varepsilon$, rejecting $H_{0}:P\in\mathcal{P}_{0}$ whenver $M_{n}$ exceeds $\varepsilon$. We refer to test (super)martingales, e-processes, and ENSMs for $\mathcal{P}_{0}$ as test processes for the null $\mathcal{P}_{0}$.
A confidence sequence for a parameter $\theta$ follows in the same manner as mentioned in Section 2.2 if we have an NSM or an ENSM $\{M^{\theta_{0}}_{n}\}$ for every $\mathcal{P}_{\theta_{0}}=\{P\in\mathcal{P}:\theta(P)=\theta_{0}\}$: by solving the inequality $M_{n}^{\theta_{0}}\leqslant 1/\alpha$ or $M_{n}^{\theta_{0}}\leqslant\varepsilon$ in $\theta_{0}$, that is. An illustrative comparison between NSMs and ENSMs in the Gaussian case with known variance for both constructing tests and confidence sequences is provided recently by Wang and Ramdas [2023a, Section 5].
@@
2.4 e-Power of Test Processes
Desirably, a powerful test process shrinks or stays small under the null $\mathcal{P}_{0}$, and grows under the alternative $\mathcal{P}_{1}$ so rejection can happen. To measure the power of a test process under a fixed alternative distribution $Q\in\mathcal{P}_{1}$, we define the e-power of test processes as follows [Vovk and Wang, 2024].
Definition 2.3 (e-power).
For a test process $\{M_{n}\}$, we define its e-power as the $Q$-almost sure limit $\liminf_{n\to\infty}\frac{\log M_{n}}{n}$. For a single e-value $M=M(X_{1},\dots,X_{n})$, we define its e-power as $\mathrm{E}_{Q}(\log M)$.
The motivation, we note, is that it quantifies the rate at which the process grows exponentially under the alternative, i.e., how fast evidence accumulates. The value of an e-process is often intuitively understood as one’s total capital when betting against a casino pricing according to the null (hence the term “game-theoretic statistics”, see e.g. Shafer and Vovk [2005]). The idea of quantifying and optimizing the growth of a capital process by its expected log-value can be traced back to Kelly [1956], Breiman [1961], among other classical references in the literature. Breiman [1961], for example, justifies this log-optimality criterion by noting that “under reasonable betting systems $M_{n}$ increases exponentially and maximizing $\mathrm{E}_{Q}(\log M_{n})$ maximizes the rate of growth.”
More concretely, many e-processes in the literature take a multiplicative form $M_{n}=\prod_{i=1}^{n}M(X_{i})$ where $M(X_{i})$’s are i.i.d. e-values. In this case, the e-power of the entire process $\{M_{n}\}$ and that of the individual e-values $M(X_{i})$ coincide due to the strong law of large numbers, and the growth is characterized asymptotically as $M_{n}=\exp(n\mathrm{E}_{Q}(\log M_{1})+o_{Q}(n))$. Alternatively, one can also informally say that the e-power $\mathrm{E}_{Q}(\log M_{1})$ is roughly inversely proportional to the expectation of the rejection time $\tau=\min\{n:M_{n}\geqslant 1/\alpha\}$ as
$\log(1/\alpha)\approx\mathrm{E}_{Q}(\log M_{\tau})=\mathrm{E}_{Q}(\log M_{1})\cdot\mathrm{E}_{Q}(\tau),$ (4)
where the equality is due to Wald’s identity (see e.g. Roters [1994, Theorem 2.1]). More formally, lower bounds are known for the expected rejection time of simple testing problems, attained exactly by the inverse of e-power of the likelihood ratio martingale, a topic we shall return to in Section 2.6.